# Interleaved dual-species arrays of single atoms using a passive optical element and one trapping laser

### v2

@@
Abstract
We demonstrate trapping of individual rubidium (Rb) and cesium (Cs) atoms in an interleaved array of bright tweezers and dark bottle-beam traps, using a microfabricated optical element illuminated by a single laser beam. This trapping approach is simple and robust compared to common approaches that use active optoelectronic components and multiple laser wavelengths, and can be easily adapted for other atomic species.
@@ 1 Introduction
Arrays of trapped neutral atoms are being rapidly developed for quantum sensing, communication, and computing applications[1]. Scaling up to large numbers of trapped atoms in well-ordered arrays is necessary for large-scale computation and other quantum-information applications. Large arrays have been achieved using sophisticated optical systems comprising high-power lasers and active beam-shaping components such as acousto-optic deflectors (AODs) [2], spatial light modulators (SLMs) [3, 4], and digital mirror devices (DMDs) [5, 6]. New developments utilize interleaved arrays of two different atomic species for error correction and noise reduction[7, 8, 9, 10] where, for example, one type of atom can be a computation qubit and the other a measurement qubit. Trapping multiple species of atoms simultaneously adds to system complexity, especially when using multiple trapping lasers with different wavelengths and active electro-optical components [8]. System complexity can be reduced by using specially fabricated passive components to generate arrays of optical traps[11, 12, 13, 14, 15].
@@ 2 Working principle and design
To enable simultaneous two-species trapping of 87Rb and 133Cs at different sites with a single laser, we select a laser wavelength between the Rb D1 line (795 nm) and the Cs D2 line (852 nm), such that the dynamic polarizabilities of Rb and Cs have opposite signs [Fig. 1(b)]. In this scheme, a region of high intensity is attractive to Rb atoms, but repulsive to Cs atoms, and vice versa. We note that there are additional windows of opposing polarizabilities in the nm and nm ranges, however these regions fall near atomic resonances, resulting in high scattering rate and significantly mismatched polarizabilities, potentially causing unequal trap depths between the two species.
@@
In our 4 imaging system, the spatial filter is placed at the back focal plane of the first lens (i.e., in the Fourier plane). The amplitude distribution at the filter plane is the Fourier transform of the amplitude distribution at the mask (flat-top beams for bright traps, and flat-bottom beams for dark traps). The spatial filter is used to remove the high-frequency components to make the trapping profile at the image plane Gaussian-like, which localizes the atoms at the centers of the traps. To achieve the best quality of the traps, certain conditions need to be satisfied, especially the size relationships between the radius of the apertures , the radius of the opaque disks , and the iris radius [Fig. 1(a)].
In the case of bright traps, the beam-amplitude profile after the first lens with focal lens becomes a Fourier transform of the flat-top beam, which is , i.e., an airy disk ( is the first-order Bessel function of the first kind, and is the radial distance from the center of each disk). To maximize the depth of the bright traps with a single aperture, the iris radius needs to block the light past the radius where the amplitude goes from positive to zero for the first time, which is shown in Eq. (1)
@@
As discussed previously, both bright and dark traps share a single iris in the Fourier plane. Thus the iris radii in eqs. (1) and (2) are equal, which gives the necessary relationship between the the dark and bright trap radii:
@@ 3 Materials and non-ideal phases
To achieve the desired transmittance in the background region, we used a thin layer of material with an intermediate degree of absorption such that (a) a subwavelength-thick layer can result in enough absorption of the incident light without a significant phase shift, and (b) the absorption is not so high, such that the transmittance can be controlled on the single-percent level by changing deposition conditions such as deposition time. We found that amorphous semiconductors with band gaps smaller than the photon energy, such as amorphous germanium (a-Ge), can work in this role. Here, we chose electron-beam evaporated Ge, which can be smooth for thin layers [20]. We used the transfer-matrix method (TMM) to calculate the transmittance and phase shift in the background region coated with a-Ge, and the results are shown in Fig. 3(b, c).
Based on the calculated transmittance and phase profiles, we performed a simulation consisting of a Fourier transform of the electric field after the mask to yield the field distribution in the Fourier plane, spatial filtering at the Fourier plane, then an inverse Fourier transform to find the field distribution at the image plane. This simulation considers interference between the dark and bright traps, which enables us to precisely calculate what a-Ge thickness is required to achieve the desired transmittance to satisfy Eq. (4), shown in Fig. 2(c). When the background material thickness is not optimal, this mask can still be used to trap both species of atoms, but requires using a higher laser power.
@@ 5 Trapping Rb and Cs atoms
In our experiment, the mask was illuminated with an 825 nm diode laser amplified by a tapered amplifier (MOGLabs). The incident beam was diffracted using an acousto-optic modulator for intensity control and switching, and then coupled into a single-mode optical fiber. To produce more uniform traps, we used a diffractive top-hat beam generator to generate a square uniform beam profile. This minimizes the spatial variation of trap depths across the array compared to a Gaussian incident beam. After the intensity mask, the beam was de-magnified by a 4 optical system with mm and mm and shaped by a variable iris on an x-y translation stage mounted in the Fourier plane. The optical setup is shown in Fig. 5(a). A CCD camera is placed at the focal plane of lens 2 to monitor the shape of the trapping beams before further de-magnification into the vacuum cell. The iris was centered on the beam by closing it all the way and optimizing the x-y symmetry of the resulting pattern. Then the iris was opened and adjusted while monitoring the image to maximize the brightness of the bright traps. Fig. 5(b) shows an image of the trap array on the CCD camera. Fig. 5(c) shows the Gaussian-like line profiles of the dark and bright traps highlighted in Fig. 5(b) near the center of the array. Due to higer-than-ideal transmittance in the background region, the trap depths of the dark traps are deeper than the bright traps, but this can be addressed by adjusting the background transmittance level.
@@
After the intermediate-step imaging, a mirror was inserted in the beam path, then a sequence of lenses with focal lengths was used to relay the transmitted light into the vacuum cell for atom trapping (shown in Fig. 6(a)). The mm lens is a custom objective lens with high numerical aperture (NA) and correction for the vacuum cell wall, suitable for creating a diffraction-limited image of the trap array at the center of a glass vacuum cell where 87Rb and 133Cs were collected in a dual-species 3D magneto-optical trap (MOT). The de-magnification from the mask to the atoms was , corresponding to a periodicity of m with a total power of 450 mW. Considering the input laser power is approximately 650 mW, the power efficiency of this specific imaging system is about .
@@
Atoms were first loaded into a dual-species MOT. The magnetic quadrupole gradient was abruptly turned off and the detuning of the cooling light from resonance was increased to further cool the atoms via polarization gradient cooling (PGC) [23]. During the PGC phase, atoms were stochastically loaded into the optical tweezer (bright) and bottle (dark) traps, where they underwent collisional blockade [24] until at most one atom per trap site survived. The resulting atom arrays were then imaged by pulsing quasi-resonant light from the side of the vacuum cell (shown in Fig. 6(a)) to induce fluorescence at wavelengths 780 nm for Rb and 852 nm for Cs. This fluorescence was collected via a second high-NA objective on the other side of the vacuum cell. A series of dichroic beamsplitters was used to separate the trapping light from the fluorescence and to split the fluorescence light into two separate imaging paths, one for each species, before recombining the images of the two interleaved arrays onto a single electron-multiplying charge-coupled device (EMCCD, Nuvu HNü512). The imaging separation of Rb and Cs atoms helps increase the signal to background of each species. The averaged fluorescence image of 5000 loading shots collected by the EMCCD is shown in Fig. 6 (a) with the Rb array of atoms in the lower-left and the Cs array in the upper-right. Each atom-trapping site is spatially separated by about 6 .
@@
In Fig. 6 (b), we reconstructed the atomic fluorescence images of Cs and Rb by aligning the centers of each array. As per the design, the two arrays of different species are interleaved. This image is reconstructed by defining the intensity values of Cs as positive and the intensity values of Rb as negative, then summing them. Cs atoms are shown as yellow [RGB = (1, 1, 0)], and Rb atoms are shown as blue [RGB = (0, 0, 1)]. The line profiles of the fluorescence from Cs and Rb atoms are shown in Fig. 6 (c). In Fig. 6 (d), the atomic fluorescence histograms of Cs and Rb at selected sites are plotted. In each histogram, the left peak is from background noise and the right peak indicates single atom occupation of the trap.
@@ 6 Conclusion
We demonstrated simultaneous trapping of individual Rb and Cs atoms, positioned in an atom-specific interleaved grid, using a single trapping laser and an optical system comprising a three-transmittance-level optical mask and a spatial filter. This is a simpler approach to trapping arrays of multiple atomic species compared to the use of active optical components like SLMs, AODs, and DMDs, and multiple lasers. The array generator works for a relatively wide range of wavelengths and is robust to fabrication imperfections. The approach can be readily adapted to any desired configuration of atom traps at any wavelength, and will therefore be useful for a variety of single-atom experiments and applications.
@@
7 Data availability statement
Data underlying the results is openly available in Zenodo at doi.org/10.5281/zenodo.14533096.
8 Acknowledgments

### v3

@@
Interleaved dual-species arrays of single atoms using a passive optical element and one trapping laser
[…]
We demonstrate trapping of individual rubidium (Rb) and cesium (Cs) atoms in an interleaved array of bright tweezers and dark bottle-beam traps, using a microfabricated optical element illuminated by a single laser beam and a 4F system with spatial filtering. Our approach exploits the opposite-sign dynamic polarizabilities of Rb and Cs, ensuring each species is exclusively trapped in either bright or dark sites. The passive optical mask creates optimal trap depths for both species using three transmittance levels while minimizing the optical phase difference, implemented using a variable-thickness absorbing layer of amorphous germanium. This trapping architecture achieves atom loading rates close to 50% while reducing system complexity compared to conventional methods using active optoelectronic components and/or multiple laser wavelengths.
@@ 1 Introduction
Arrays of trapped neutral atoms are being rapidly developed for quantum sensing, communication, and computing applications[1]. Scaling up to large numbers of trapped atoms in well-ordered arrays is necessary for large-scale computation and other quantum-information applications. Large arrays have been achieved using sophisticated optical systems comprising high-power lasers and active beam-shaping components such as acousto-optic deflectors (AODs) [2], spatial light modulators (SLMs) [3, 4], and digital mirror devices (DMDs) [5, 6]. Recent developments utilize interleaved arrays of two different atomic species for error correction and noise reduction[7, 8, 9, 10] where, for example, one type of atom can be a computation qubit and the other a measurement qubit. Trapping multiple species of atoms simultaneously adds to system complexity, especially when using multiple trapping lasers with different wavelengths and active electro-optical components [8]. System complexity can be reduced by using specially fabricated passive components to generate arrays of optical traps[11, 12, 13, 14, 15, 16].
@@ 2 Working principle and design
To enable simultaneous two-species trapping of 87Rb and 133Cs at different sites with a single laser, we select a laser wavelength between the Rb D1 line (795 nm) and the Cs D2 line (852 nm), such that the dynamic polarizabilities of Rb and Cs have opposite signs [Fig. 1(b)]. In this scheme, a region of high intensity is attractive to Rb atoms, but repulsive to Cs atoms, and vice versa. We note that there are additional windows of opposing polarizabilities in the nm and nm ranges, however these regions fall near atomic resonances, resulting in high scattering rate and substantially mismatched polarizabilities, potentially causing unequal trap depths between the two species.
@@
In our 4 imaging system (Fig. 1(a)), the spatial filter is placed at the back focal plane of the first lens (i.e., in the Fourier plane). The amplitude distribution after the optical mask is an array of flat-top/bottom beams, and can be understood as the convolution between the electric field of a single flat-top/bottom (bright/dark) beam with a finite Dirac comb [Fig. 2(a)]. For simplicity, the background amplitude is not discussed here since it is a DC signal which is not filtered by the aperture in the Fourier plane. At the Fourier plane after the first lens (Fig. 2(b)), the electric field can be understood via the Convolution Theorem to be the product of the Fourier transform of the flat-top/bottom beam, and the Fourier transform of the finite Dirac comb. In this plane, we place an iris (gray area in Fig. 2(b)), which is a spatial filter that removes the high-frequency components. At the image plane (Fig. 2(c)), the field profile is the convolution of the inverse Fourier transform of each function in Fig. 2(b)). Note that the spatial filter actually also applies to the Fourier transform of the Dirac comb in Fig. 2(b), but whether or not you apply it is irrelevant because those components end up being multiplied by 0.
The spatial filter is used to remove the high-frequency components to make the trapping profile at the image plane Gaussian-like, which localizes the atoms at the centers of the traps. To achieve the best quality of the traps, certain conditions need to be satisfied, especially the size relationships between the radius of the apertures , the radius of the opaque disks , and the iris radius [Fig. 1(a)].
In the case of bright traps, the beam-amplitude profile after the first lens with focal length becomes a Fourier transform of the flat-top beam, which is , i.e., an airy disk (cross-section shown in Fig. 2 (b), is the first-order Bessel function of the first kind, and is the radial distance from the center of each disk). To maximize the depth of the bright traps with a single aperture, the iris radius needs to block the light past the radius where the amplitude goes from positive to zero for the first time, which is shown in Fig. 2(b, left) and Eq. (1):
@@
As discussed previously, both bright and dark traps share a single iris in the Fourier plane (Fig. 2(b)). Thus the iris radii in eqs. (1) and (2) are equal, which gives the necessary relationship between the dark and bright trap radii:
@@ 3 Materials and non-ideal phases
To achieve the desired transmittance in the background region, we used a thin layer of material with an intermediate degree of absorption such that (a) a subwavelength-thick layer can result in enough absorption of the incident light without too much phase shift, and (b) the absorption is not so high, such that the transmittance can be controlled on the single-percent level by changing deposition conditions such as deposition time. We found that amorphous semiconductors with band gaps smaller than the photon energy, such as amorphous germanium (a-Ge), can work in this role. Here, we chose electron-beam evaporated Ge, which can be smooth for thin layers [21]. We used the transfer-matrix method (TMM) to calculate the transmittance and phase shift in the background region coated with a-Ge, and the results are shown in Fig. 4(b, c). The wavelength-dependent refractive index and extinction coefficient are adapted from Ref. [22].
Based on the calculated relative transmittance and phase difference between the background the transparent region, we performed a simulation similar to Fig. 2, consisting of a Fourier transform of the electric field after the mask to yield the field distribution in the Fourier plane, spatial filtering at the Fourier plane, then an inverse Fourier transform to find the field distribution at the image plane. In the previous simulation, zero phase difference between the background and the transparent region was assumed. Now we include the accumulated phase in the transparent region, , and in the background region, , and the resulting complex-field spatial distribution is then fed into the calculation described in Fig. 2. This simulation considers interference between the dark and bright traps, which enables us to precisely calculate what a-Ge thickness is required. When the background material thickness is not optimal, this mask can still be used to trap both species of atoms, but requires using a higher laser power.
@@ 5 Trapping Rb and Cs atoms
In our experiment, the mask was illuminated with an 825 nm diode laser amplified by a tapered amplifier (MOGLabs). The incident beam was diffracted using an acousto-optic modulator for intensity control and switching, and then coupled into a single-mode optical fiber. After the intensity mask, the beam was de-magnified by a 4 optical system with mm and mm and shaped by a variable iris on an x-y translation stage mounted in the Fourier plane. The optical setup is shown in Fig. 6(a), where a top-hat generator was also used to form a uniform intensity profile. A CCD camera is placed at the focal plane of lens 2 to monitor the shape of the trapping beams before further de-magnification into the vacuum cell. The iris was centered on the beam by closing it all the way and optimizing the x-y symmetry of the resulting pattern. Then the iris was opened and adjusted while monitoring the image to maximize the brightness of the bright traps. Fig. 6(b) shows an image of the trap array on the CCD camera. Fig. 6(c) shows the Gaussian-like line profiles of the dark and bright traps highlighted in Fig. 6(b) near the center of the array. Due to higer-than-ideal transmittance in the background region, the trap depths of the dark traps are deeper than the bright traps, but this can be addressed by adjusting the background transmittance level.
@@
After the intermediate-step imaging, a mirror was inserted in the beam path, then a sequence of lenses with focal lengths was used to relay the transmitted light into the vacuum cell for atom trapping (shown in Fig. 7(a)). However, the top-hat beam generator is removed before further experiment for simplicity. The mm lens is a custom objective lens with high numerical aperture (NA) and correction for the vacuum cell wall, suitable for creating a diffraction-limited image of the trap array at the center of a glass vacuum cell where 87Rb and 133Cs were collected in a dual-species 3D magneto-optical trap (MOT). The de-magnification from the mask to the atoms was , corresponding to a periodicity of m with a total power of 450 mW. Considering the input laser power is approximately 650 mW, the power efficiency of this specific imaging system is about .
@@
Atoms were first loaded into a dual-species MOT. The magnetic quadrupole gradient was abruptly turned off and the detuning of the cooling light from resonance was increased to further cool the atoms via polarization gradient cooling (PGC) [25]. During the PGC phase, atoms were stochastically loaded into the optical tweezer (bright) and bottle (dark) traps, where they underwent collisional blockade [26] until at most one atom per trap site survived. The resulting atom arrays were then imaged by pulsing quasi-resonant light from the side of the vacuum cell (shown in Fig. 7(a)) to induce fluorescence at wavelengths 780 nm for Rb and 852 nm for Cs. This fluorescence was collected via a second high-NA objective on the other side of the vacuum cell. A series of dichroic beamsplitters was used to separate the trapping light from the fluorescence and to split the fluorescence light into two separate imaging paths, one for each species, before recombining the images of the two interleaved arrays onto a single electron-multiplying charge-coupled device (EMCCD, Nuvu HNü512). The imaging separation of Rb and Cs atoms helps increase the signal to background of each species. The averaged fluorescence image of 5000 loading shots collected by the EMCCD is shown in Fig. 7 (a) with the Rb array of atoms in the lower-left and the Cs array in the upper-right. The exposure time of each fluorescence measurement is 50 ms. Each atom-trapping site is spatially separated by about 6 . The atom lifetime in the traps was observed to be seconds to tens of seconds.
In the averaged fluorescence image, both Cs and Rb arrays have background some fluorescence around them. We understand this background to consist of three components: (1) A large portion is due to scattering the 852 nm and 780 nm probe light off of the cell that makes it to the camera. We observed that when the MOT is off, the EMCCD still collects substantial light from the probe lasers. Also, when we block part of the probe laser, shadows show up on the EMCCD; (2) additional background is due the atoms trapped at the Talbot planes (Fig. S6). When we scanned the objective lens in the z direction, we observed atoms trapped in planes other than the image plane; (3) for the Rb array, some background light is caused by the fluorescence of the atoms being trapped in the Gaussian envelope (Fig. S7(a, b)) created by the input Gaussian beam. The Gaussian envelope behaves as a much larger bright trap. This background noise can be mitigated by using a top-hat beam instead of Gaussian beam. As a result, the fluorescence-based readout of Rb atoms generally has lower readout fidelity in our trapping experiment, which we will discuss below.
@@
In Fig. 7(b), we reconstructed the atomic fluorescence images of Cs and Rb by aligning the centers of each array. As per the design, the two arrays of different species are interleaved. This image is displayed by defining the intensity values of Cs as positive and the intensity values of Rb as negative, then summing them. Cs atoms are shown as yellow [RGB = (1, 1, 0)], and Rb atoms are shown as blue [RGB = (0, 0, 1)]. The line profiles of the fluorescence from Cs and Rb atoms are shown in Fig. 7(c). In Fig. 7(d), the atomic fluorescence histograms of Cs and Rb at selected sites are plotted. The fluorescence histogram is fitted with a sum of two Gaussian distributions, one centered at the background signal intensity level and the other centered at a higher intensity level, indicating a single atom. Once the double-Gaussian model has been fitted, the intensity at the lowest probability density level between the two peaks is set as the threshold intensity (dashed lines in Fig. 7(d)). When the intensity is higher than the threshold, the trap is considered to be occupied, and vice versa. Based on the fitting parameters and the threshold intensity, we calculated the readout fidelity [27], which is listed on top of Fig. 7(d). For the entire array of both Rb and Cs atoms, the fitting of the histograms and the calculated readout fidelities for each trapping site are plotted in supplementary Fig. S3 and Fig. S4.
@@ 6 Conclusion
We demonstrated simultaneous trapping of individual Rb and Cs atoms, positioned in an atom-specific interleaved grid, using a single trapping laser and an optical system comprising a three-transmittance-level optical mask and a spatial filter. This is a simpler approach to trapping arrays of multiple atomic species compared to the use of active optical components like SLMs, AODs, and DMDs, and multiple lasers. The array generator works for a relatively wide range of wavelengths, is robust to fabrication imperfections, and can be scaled to arbitrarily large array sizes by scaling the laser power with the number of desired traps. The approach can be readily adapted to any desired configuration of atom traps at any wavelength, and will therefore be useful for a variety of single-atom experiments and applications.
@@
Competing Interests
CF, SD, PH, MAK, and MS are inventors on a patent application (P240121US01) submitted by the Wisconsin Alumni Research Foundation (WARF), covering devices and methods to trap arrays of isolated particles of multiple species. All other authors declare they have no competing interests.
Funding
@@
Author Contributions
Conceptualization: CF, MAK, PH, MS. Calculations and simulations: CF, SD, SAN, PH. Experimental realization: CF, JM, JG, ML, MG, MB, SAN. Data analysis: CF, JM, JG, MAK, MS Project administration: MAK, MS. Writing original draft: CF, SAN, MAK, MS. Review and editing: all authors.
Data and Materials Availability
Data underlying the results is openly available in a Zenodo repository at doi.org/10.5281/zenodo.14533096.
Supplementary Materials
Ge thin film transmittance stability
To evaluate the long-term stability of the evaporated Ge thin film’s optical properties, especially the transmittance, we measured the relative transmittance of the Ge-coated background region versus the transparent region after four months of using the optical mask. The optical mask is used in an atmospheric environment with 808 nm laser exposure at around 1 W. As shown in Fig. S1, no substantial changes in transmittance are observed, indicating that the Ge thin film maintains its optical properties under high-power laser exposure.
Escape position of the traps
Cs array histogram
The fluorescence-based readout of Cs atoms at each trapping site across the entire array is shown in Fig.S3. Double-Gaussian distribution model is used to fit the histograms. The fitting parameters are used to determine the threshold intensity (vertical dashed line in Fig.S3), then calculate the loading rate and the readout fidelity.
Rb array histogram
The fluorescence-based readout of Rb atoms at each trapping site across the entire array is shown in Fig.S4. The fitting parameters are used to determine the threshold intensity (vertical dashed line in Fig.S4), then calculate the loading rate and the readout fidelity.
Imaging the dual-species array with less coma aberration
In Fig. 7(a), the averaged fluorescence image of the Rb array is distorted. This is because the imaging path of Rb array is slightly off axis. In the fluorescence readout process, we spatially separate the Cs and Rb arrays using dichroic beam splitters. However, this introduces coma aberration on the Rb array. To verify that the distortion can be fixed by better optical alignment or a reduction of the aberrations, another set of data was collected (Fig. S5 with 2000 loading shots. The Rb array shows less coma aberration.
Talbot effect
Since the trapping profile has periodic structure, the Talbot effect causes self-images of the intensity pattern at the light propagation direction, shown in Fig. S6. This unintended self-imaging might blur fluorescence images. In our trapping experiment, we have noticed the atom loading at Talbot planes when the objective lens is scanned in Z direction.
Intensity profile with Gaussian input and resulting light shift
In Fig. S7(a), we calculated the intensity profile in the image plane, where total power is around 450 mW. Fig. S7(b) shows the line profiles through the array for dark traps (blue) and bright traps (red). For the Rb atoms in the bright traps, the intensities of the trap centers change substantially over the array. For the Cs atoms in the dark traps, the intensities at the trapping sites are all close to zero. The light shift is proportional to the laser intensity at the trapping site, so the light shift of the Rb array is less homogeneous than the Cs array when Gaussian beam is used on the intensity mask. A top-hat beam generator can greatly improve the homogeneity of the bright trap array.